# A Hierarchy of Spectral Gap Certificates for Frustration-Free Spin Systems

### 06 Nov 2024 v1

@@ Appendix C Proof for inclusion of finite-size based methods
where the positivity follows because the RHS is sum of disjoint anticommutators. Thus, we explicitly found a $n-$local positive operator decomposition for $[H^{2}]_{n}-\Delta_{\mathrm{Knabe}}(n)$, proving that $\Delta_{\mathrm{Knabe}}(n)$ lies in the feasible set of the LTI SDP.
C.2 Gosset–Mozgunov bound eq. 23
Similar to the previous section, we will derive the positive $n-$local term $q_{n}(\Delta_{\mathrm{GM}}(n))$ for the Gosset–Mozgunov bound which has the expression
$\Delta_{m}^{\circ}\geq\Delta_{\mathrm{GM}}(n)=\frac{5}{6}\frac{n^{2}+n}{n^{2}-4}\left(\epsilon_{n}-\frac{6}{n^{2}+n}\right),$ (95)
where it lower bounds the gap of system size $m>2n$. Expanding the global operator $Q_{n}(\Delta_{\mathrm{GM}}(n))$, we obtain
$Q_{n}(\Delta_{\mathrm{GM}}(n))$ $=H+\{h_{i},h_{i+1}\}+...+\{h_{i},h_{i+n-2}\}-\Delta_{\mathrm{GM}}(n)H,$ (96)
$=(1-\Delta_{\mathrm{GM}}(n))H+\sum_{i}\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\}$ (97)
Now, let us simplify the coefficient $(1-\Delta_{\mathrm{GM}}(n))$ of the first term as,
$1-\Delta_{\mathrm{GM}}(n)=a-b\epsilon_{n},$ (98)
where we defined
$a=\frac{n^{2}+1}{n^{2}-4},\quad b=\frac{5}{6}\frac{n^{2}+n}{n^{2}-4}$ (99)
for better readability of the upcoming steps. In eq. 97,
$Q_{n}(\Delta_{\mathrm{GM}}(n))$ $=\sum_{i}(ah_{i}-b\epsilon_{n}h_{i})+\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\}$ (100)
The following discussion heavily uses notation from Ref. [20]. Firstly, define
$B_{n,i}=\sum_{j=0}^{n-2}c_{j}h_{i+j}$ (101)
where coefficients $c_{j}$ satisfy the relations from equations (5)-(7) of Ref. [20]. Then, expanding $B_{n,i}^{2}$, we obtain
$B_{n,i}^{2}$ $=\sum_{j=0}^{n-2}c_{j}^{2}h_{i+j}+\sum_{j=0}^{n-3}c_{j}c_{j+1}\{h_{i+j},h_{i+j+1}\}+\sum\limits_{j,k\in\{0,..,n-2\};|j-k|\geq 2}c_{j}c_{k}h_{i+j}h_{i+k}.$ (102)
Using the LTI freedom (App. A), we can perform the following replacement of $h_{i}$ in the first term eq. 100 with
$h_{i}\to\frac{1}{\sum_{r=0}^{n-2}c_{r}^{2}}\sum_{j=0}^{n-2}c_{j}^{2}h_{i+j}$ (103)
and the $h_{i}$ in the second term of eq. 100, with
$h_{i}$ $\to\frac{B_{n,i}}{\sum_{r=0}^{n-2}c_{r}}.$ (104)
Note that as for the previous proof for the Knabe bound, we can perform these replacements (using LTI condition) because they do not modify the complete sum over the index $i$. Substituting the above two equations into eq. 100,
$Q_{n}(\Delta_{\mathrm{GM}}(n))$ $=\sum_{i}\left(\sum_{j=0}^{n-2}\frac{ac_{j}^{2}}{\sum_{r=0}^{n-2}c_{r}^{2}}h_{i+j}-b\epsilon_{n}\frac{B_{n,i}}{\sum_{r=0}^{n-2}c_{r}}\right)+\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\}.$ (105)
The next step is to use the LTI freedom to change the local form of the anticommutators,
$\{h_{i},h_{i+1}\}$ $\to\frac{\sum_{j=0}^{n-3}c_{j}c_{j+1}\{h_{i+j},h_{i+j+1}\}}{\sum_{r=0}^{n-3}c_{r}c_{r+1}}$ (106)
$\sum_{x=2}^{n-2}\{h_{i},h_{i+x}\}$ $\to\frac{\sum\limits_{j,k\in\{0,..,n-2\};|j-k|\geq 2}c_{j}c_{k}h_{i+j}h_{i+k}}{\sum\limits_{q,r\in\{0,..,n-2\};|q-r|\geq 2}c_{q}c_{r}}.$ (107)
keeping the global sum of eq. 100 the same. Substituting the above two equations into eq. 105,
$\begin{split}Q_{n}(\Delta_{\mathrm{GM}}(n))=\sum_{i}\left(\sum_{j=0}^{n-2}\frac{ac_{j}^{2}}{\sum_{r=0}^{n-2}c_{r}^{2}}h_{i+j}-b\epsilon_{n}\frac{B_{n,i}}{\sum_{r=0}^{n-2}c_{r}}\right)&+\frac{\sum_{j=0}^{n-3}c_{j}c_{j+1}\{h_{i+j},h_{i+j+1}\}}{\sum_{r=0}^{n-3}c_{r}c_{r+1}}\\ &+\frac{\sum\limits_{j,k\in\{0,..,n-2\};|j-k|\geq 2}c_{j}c_{k}h_{i+j}h_{i+k}}{\sum\limits_{q,r\in\{0,..,n-2\};|q-r|\geq 2}c_{q}c_{r}}.\end{split}$ (108)
Now, using the above expansion, we define $q_{n}(\Delta_{\mathrm{GM}}(n))$ as,
$\begin{split}q_{n}(\Delta_{\mathrm{GM}}(n)):=\left(\sum_{j=0}^{n-2}\frac{ac_{j}^{2}}{\sum_{r=0}^{n-2}c_{r}^{2}}h_{i+j}-b\epsilon_{n}\frac{B_{n,i}}{\sum_{r=0}^{n-2}c_{r}}\right)&+\frac{\sum_{j=0}^{n-3}c_{j}c_{j+1}\{h_{i+j},h_{i+j+1}\}}{\sum_{r=0}^{n-3}c_{r}c_{r+1}}\\ &+\frac{\sum\limits_{j,k\in\{0,..,n-2\};|j-k|\geq 2}c_{j}c_{k}h_{i+j}h_{i+k}}{\sum\limits_{q,r\in\{0,..,n-2\};|q-r|\geq 2}c_{q}c_{r}}.\end{split}$ (109)
And the requirement is to prove the positivity for the above choice of
$q_{n}(\Delta_{\mathrm{GM}}(n))\overset{!}{\succeq}0.$ (110)
The first step is to use the result of Lemma 4 of Ref. [20], which is the relation
$B_{n,i}^{2}\succeq\epsilon_{n}\left(\frac{1}{n-1}\sum_{j=0}^{n-2}c_{j}\right)B_{n,i}.$ (111)
Secondly, we replace the $B_{n,i}$ in the second term of eq. 109 with $B_{n,i}^{2}$ using the above relation,
$q_{n}(\Delta_{\mathrm{GM}}(n))$ $\succeq\frac{a\sum_{j=0}^{n-2}c_{j}^{2}h_{i+j}}{\sum_{r=0}^{n-2}c_{r}^{2}}-\frac{b(n-1)B_{n,i}^{2}}{\left(\sum_{r=0}^{n-2}c_{r}\right)^{2}}+\frac{\sum_{j=0}^{n-3}c_{j}c_{j+1}\{h_{i+j},h_{i+j+1}\}}{\sum_{r=0}^{n-3}c_{r}c_{r+1}}+\frac{\sum\limits_{j,k\in\{0,..,n-2\};|j-k|\geq 2}c_{j}c_{k}h_{i+j}h_{i+k}}{\sum\limits_{q,r\in\{0,..,n-2\};|q-r|\geq 2}c_{q}c_{r}}$ (112)
$=\sum_{j=0}^{n-2}\left(\frac{a}{\sum_{r=0}^{n-2}c_{r}^{2}}-\frac{b(n-1)}{\left(\sum_{r=0}^{n-2}c_{r}\right)^{2}}\right)c_{j}^{2}h_{i+j}+\left(\frac{1}{\sum_{r=0}^{n-3}c_{r}c_{r+1}}-\frac{b(n-1)}{\left(\sum_{r=0}^{n-2}c_{r}\right)^{2}}\right)\sum_{j=0}^{n-3}c_{j}c_{j+1}\{h_{i+j},h_{i+j+1}\}$ (113)
$\quad\quad\quad\quad\quad\quad\quad+\sum\limits_{j,k\in\{0,..,n-2\};|j-k|\geq 2}\left(\frac{1}{\sum\limits_{q,r\in\{0,..,n-2\};|q-r|\geq 2}c_{q}c_{r}}-\frac{b(n-1)}{\left(\sum_{r=0}^{n-2}c_{r}\right)^{2}}\right)c_{j}c_{k}h_{i+j}h_{i+k}$
There are three terms based on $c_{j}$’s, that can be simplified using relations from equations (30)-(32) of Ref. [20], which are
$\sum_{r=0}^{n-2}c_{r}$ $=\frac{n^{3}-n}{6}$ (114)
$\sum_{r=0}^{n-2}c_{r}^{2}$ $=\frac{n^{5}-n}{30}$ (115)
$\sum_{r=0}^{n-3}c_{r}c_{r+1}$ $=\frac{n^{5}-5n^{3}+4n}{30}.$ (116)
Substituting the above relations into eq. 113, we obtain that the coefficient of second term is exactly zero, i.e.
$\left(\frac{1}{\sum_{j=0}^{n-3}c_{j}c_{j+1}}-\frac{b(n-1)}{\left(\sum_{j=0}^{n-2}c_{j}\right)^{2}}\right)=0.$ (117)
Similarly, upon simplifying the coefficient of the first term,
$\left(\frac{a}{\sum_{r=0}^{n-2}c_{r}^{2}}-\frac{b(n-1)}{\left(\sum_{r=0}^{n-2}c_{r}\right)^{2}}\right)=30\left(\frac{(n^{2}+1)}{n(n^{4}-1)(n^{2}-4)}-\frac{n^{2}+1}{n(n^{4}-1)(n^{2}-4)}\right)=0,$ (118)
which is also zero. Finally, the coefficient of the last term is simplified using the autocorrelation lemma from Ref. [20],
$\sum\limits_{j,k\in\{0,..,n-2\};|j-k|\geq 2}c_{j}c_{k}\leq\sum_{j=0}^{n-3}c_{j}c_{j+1}.$ (119)
Substituting it in the last term of eq. 113, we find
$\left(\frac{1}{\sum\limits_{j,k\in\{0,..,n-2\};|j-k|\geq 2}c_{j}c_{k}}-\frac{b(n-1)}{\left(\sum_{j=0}^{n-2}c_{j}\right)^{2}}\right)\geq\left(\frac{1}{\sum_{j=0}^{n-3}c_{j}c_{j+1}}-\frac{b(n-1)}{\left(\sum_{j=0}^{n-2}c_{j}\right)^{2}}\right)=0.$ (120)
Combining all three terms, we finally obtain
$q_{n}(\Delta_{\mathrm{GM}}(n))\succeq 0.$ (121)
Thus, we explicitly derived a $n-$local positive operator decomposition for $[H^{2}]_{n}-\Delta_{\mathrm{GM}}(n)$, proving that $\Delta_{\mathrm{GM}}(n)$ lies in the feasible set of the LTI SDP.
C.3 Martingale condition bound eq. 31
@@
We show that the martingale bound of the blocked Hamiltonian $K$ is contained in the feasible set of the LTI SDP. We will mainly use the inequality (Lemma 6.3(2) of [16])
$\{P_{\mathcal{I},\mathcal{I}+1},P_{\mathcal{I}+1,\mathcal{I}+2}\}\succeq-\eta(P_{\mathcal{I},\mathcal{I}+1}+P_{\mathcal{I}+1,\mathcal{I}+2}),$ (122)
The lower bound on the gap of Hamiltonian $K$ after using the above inequality is
$K^{2}\succeq(1-2\eta)K.$ (123)
We take that $K$ is supported on $n=3r$ sites of the original Hamiltonian. Therefore, to prove that the above bound is included in the feasible set, we find a decomposition for $Q_{n}(1-2\eta)$ such that
$q_{n}(1-2\eta)\overset{!}{\succeq}0.$ (124)
We start by expanding the LHS,
$Q_{n}(1-2\eta)$ $=H+\sum_{i}(\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\})-(1-2\eta)H$ (125)
$=H+\sum_{i}(\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\})-(1-2\eta)H$ (126)
$=\sum_{i}2\eta h_{i}+(\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\})$ (127)
$=\sum_{i}\eta(h_{i}+h_{i+1})+(\{h_{i},h_{i+1}\}+\{h_{i},h_{i+2}\}+...+\{h_{i},h_{i+n-2}\})$ (128)
@@
$q_{n,i}(1-2\eta)\succeq 0.$ (130)
Appendix D Optimizing the interaction term using the SDP gradient formula
@@
First of all, we see that upon applying the SDP gradient formula ([39], Section 4.2) to the exact SDP formulation of the gap problem section II, we recover the result of the Hellmann–Feynman theorem: Let $H$ be the full system Hamiltonian with ground-state energy zero, and let $P_{0}$ be the projector to the global ground state. Assume $H$ depends on a parameter $x$ as $H(x)=H_{0}+(\mathbb{I}-P_{0})A(x)(\mathbb{I}-P_{0})$, where $A$ is some operator depending smoothly on $x$, such that $H(x)$ has the same ground state as $H_{0}$ in a neighborhood of $x_{0}$.
The SDP gradient formula tells us that the derivative of the objective function (in our case the exact gap) with respect to the parameter $x$ is given by the partial derivative of the Lagrangian at the primal-dual optimal point. In the same spirit as the Hellmann–Feynman theorem—only the explicit dependence on $x$ contributes to the gradient. Applying the SDP gradient formula to the SDP section II, where the corresponding Lagrangian is $\mathcal{L}(\rho,\delta)=\delta+\mathrm{Tr}(\rho(H^{2}-\delta H))$, and the primal-dual optimum is ($\rho^{\star}=|\psi_{1}\rangle\langle\psi_{1}|/\Delta,\delta^{\star}=\Delta$), we get
@@
Let $p_{0}$ be the projector to the ground space of $h_{0}$. Any interaction term of the form $h(X):=h_{0}+(\mathbb{I}-p_{0})X(\mathbb{I}-p_{0})$ which is still positive semidefinite and frustration free will have the same global ground state. We can use the SDP gradient formula to differentiate the gap which the SDP detects with respect to $X$. Of course we can always increase the gap arbitrarily by choosing $X=\mathrm{const.}\times\mathbb{I}$, so in addition to the positivity constraint $h(X)\geq 0$ we should impose some constraint on $X$, e.g. $||X||^{2}_{2}\leq 1$. Both of those constraints are easily differentiable and so we could perform constrained gradient optimization.
@@
Appendix E Certifiable formulation of the LTI SDP
We show the steps to ensure that lower bounds produced by the SDP do not suffer from numerical precision errors. For the LTI SDP,

### 31 Mar 2026 v2

@@ Appendix B Proof for inclusion of finite-size based methods
where the positivity follows because the RHS is sum of disjoint anticommutators. Thus, we explicitly found a $n$-local positive operator decomposition for $H^{2}]_{n}-\delta_{\mathrm{Knabe}}(n)H$, proving that $\delta_{\mathrm{Knabe}}(n)$ lies in the feasible set of the LTI SDP.
B.2 Gosset–Mozgunov bound eq. 24
Define $B_{n,k}$ as in equation (8) in Ref. [24]: $B_{n,k}=\sum_{i=k}^{n-2+k}c_{i-k}h_{i,i+1}$, where $c_{j}$ are positive coefficients satisfying the conditions stated below. Using equations (18) and (20) of Ref. [24], which relate $\sum_{k}B_{n,k}^{2}$ and $H^{2}]_{n}$, and using the expression for the Gosset–Mozgunov gap $\delta_{\mathrm{GM}}$ from equation (26) in same Ref., we obtain
$[H^{2}]_{n}-\delta_{\mathrm{GM}}H$ $=\sum_{i}q_{n,i}(\delta_{\mathrm{GM}})$ (84)
$=\frac{1}{\theta(1)}\left(\sum_{i}B_{n,i}^{2}-\frac{\sum_{j}c_{j}}{n-1}\epsilon_{n}\sum_{i}B_{n,i}+\sum_{\begin{subarray}{c}(j,k)\\ n-2\geq d(j,k)\geq 2\end{subarray}}h_{j}h_{k}\left[\theta(1)-\theta(d(j,k))\right]\right)$ (85)
where $\theta(x):=\sum_{k=0}^{n-x-2}c_{k}c_{k+x}$. The coefficients $c_{j}$ in the deformed Hamiltonian $B_{n,i}$ satisfy the following properties
  • Positive: $c_{j}>0$ for $0\leq j\leq n-2$
  • Non-decreasing upto midpoint: $c_{j}>c_{j-1}$ for $1\leq j\leq\frac{n-2}{2}$
  • Symmetric about midpoint: $c_{j}=c_{(n-2)-j}$ for $0\leq j\leq\frac{n-2}{2}$
Given the above conditions, the coefficients $c_{j}$ satisfy the 1D autocorrelation lemma [24]
$\theta(x)\geq\theta(x+1),\quad\forall\,x=0,1,...,n-3.$ (86)
This implies that the term $\theta(1)-\theta(d)$ is non negative for any $n-2\geq d\geq 2$ and thus the third term in eq. 85 is positive semidefinite.
To show inclusion of the GM bound, we want to obtain a local positive decomposition of the rest of eq. 85: $\sum_{i}B_{n,i}^{2}-\frac{\sum_{j}c_{j}}{n-1}\epsilon_{n}\sum_{i}B_{n,i}$. As in the previous section, we are free to use the LTI freedom, i.e., add a term of the form $\mathbb{I}\otimes Y-Y\otimes\mathbb{I}$. We will prove that a $Y$ exists such that $B_{n}^{2}-\frac{\sum_{j}c_{j}}{n-1}\epsilon_{n}B_{n}+\mathbb{I}\otimes Y-Y\otimes\mathbb{I}\succeq 0$ without giving an explicit expression for it (unlike in the case of the Knabe bound in the previous section).
The proof of existence of such a $Y$ relies on the following fact: For an $n$-body operator $K$ the following SDPs are dual to each other and attain the same value.
$\min_{\rho}\mathrm{Tr}K\rho$ (87)
$\text{s.t. }\rho\succeq 0$
[…]
$\max_{\epsilon,A}\epsilon$ (88)
$\text{s.t. }K+\mathbb{I}\otimes Y-Y\otimes\mathbb{I}\succeq\epsilon\mathbb{I}$
[…]
It is therefore sufficient to prove that
[…]
The proof of eq. 89 is essentially the same as lemma 4 in [24], the difference being that here $\rho$ is an $n$-body LTI state, whereas in [24] the state $\phi$ is a translation invariant eigenvector of the the $m$-site periodic chain Hamiltonian.
Proof of eq. 89: Let $G^{\perp}$ be the projector to the complement of the ground space of $B_{n}$, and let $\hat{\rho}=G^{\perp}\rho G^{\perp}/\mathrm{Tr}G^{\perp}\rho$
$\mathrm{Tr}\left[B_{n}^{2}\rho\right]=\mathrm{Tr}\left[B_{n}^{2}\hat{\rho}\right]\mathrm{Tr}G^{\perp}\rho\geq\mathrm{Tr}\left[B_{n}\hat{\rho}\right]\mathrm{Tr}\left[B_{n}\hat{\rho}\right]\mathrm{Tr}G^{\perp}\rho$ (90)
where we used the non-negativity of the variance of $B_{n}$ in $\hat{\rho}$. Due to the LTI property of $\rho$ we have:
$\mathrm{Tr}\hat{\rho}h_{i}=\frac{\mathrm{Tr}{\rho}G^{\perp}h_{i}G^{\perp}}{\mathrm{Tr}G^{\perp}\rho}=\frac{\mathrm{Tr}{\rho}h_{i}}{\mathrm{Tr}G^{\perp}\rho}=\frac{\mathrm{Tr}{\rho}h_{j}}{\mathrm{Tr}G^{\perp}\rho}=\mathrm{Tr}\hat{\rho}h_{j}$ (91)
[…]
Thus
$\mathrm{Tr}\left[B_{n}\hat{\rho}\right]=\sum_{j}c_{j}\mathrm{Tr}h_{j}\hat{\rho}=\mathrm{Tr}h\hat{\rho}\sum_{j}c_{j}=\frac{\sum_{j}c_{j}}{n-1}\sum_{i}\mathrm{Tr}h_{i}\hat{\rho}=\frac{\sum_{j}c_{j}}{n-1}\mathrm{Tr}H_{n}\hat{\rho}\geq\frac{\sum_{j}c_{j}}{n-1}\epsilon_{n}$ (92)
So we get,
$\mathrm{Tr}\left[B_{n}^{2}\rho\right]\geq\frac{\sum_{j}c_{j}}{n-1}\epsilon_{n}\mathrm{Tr}\left[B_{n}\hat{\rho}\right]\mathrm{Tr}G^{\perp}\rho=\frac{\sum_{j}c_{j}}{n-1}\epsilon_{n}\mathrm{Tr}\left[B_{n}\rho\right]$ (93)
as required.
B.3 Martingale condition bound eq. 33
@@
We show that the martingale bound of the coarse-grained projector Hamiltonian $K=\sum_{\mathcal{I}}P_{\mathcal{I},\mathcal{I}+1}$ is contained in the feasible set of the LTI SDP for gap of Hamiltonian $K$.
We modify the original $k$-local Hamiltonian into the coarse-grained projector Hamiltonian $K$ through the following steps. First, each site $\mathcal{I}$ of the coarse grained system is comprised of $l\geq k-1$ sites of the original system
$\mathcal{I}=\{(i-1)l+1,...,il\},$ (94)
And the coarse grained Hamiltonian is defined as
$\tilde{h}_{\mathcal{I},\mathcal{I}+1}=\sum\limits_{j=(\mathcal{I}-1)l+1}^{(\mathcal{I}+1)l-k+1}h_{j,...,j+k-1}.$ (95)
The locality of the new Hamiltonian $\tilde{h}_{\mathcal{I},\mathcal{I}+1}$ is $2l$-local in the original system. Next we define the corresponding projector Hamiltonian $K$,
$K=\sum_{\mathcal{I}}P_{\mathcal{I},\mathcal{I}+1},$ (96)
$\tilde{h}_{\mathcal{I},\mathcal{I}+1}\succeq\epsilon_{2l}P_{\mathcal{I},\mathcal{I}+1}.$ (97)
To prove inclusion of the martingale bound, we will mainly use the inequality (Lemma 6.3(2) of [20])
$\{P_{\mathcal{I},\mathcal{I}+1},P_{\mathcal{I}+1,\mathcal{I}+2}\}\succeq-\eta_{l}(P_{\mathcal{I},\mathcal{I}+1}+P_{\mathcal{I}+1,\mathcal{I}+2}),$ (98)
where
[…]
and $G_{i,i+1,...,i+m-1}$ denotes the projector onto the ground space of the original Hamiltonian $h$ on total $m$ sites from indices $i$ to i+m1i+m-1.
Then, $K^{2}$ can be lower bounded as
$K^{2}\geq$ $K+\sum_{\mathcal{I}}\{P_{\mathcal{I},\mathcal{I}+1},P_{\mathcal{I}+1,\mathcal{I}+2}\}$ (100)
$\geq$ $K-\eta_{l}\sum_{\mathcal{I}}(P_{\mathcal{I},\mathcal{I}+1}+P_{\mathcal{I}+1,\mathcal{I}+2})=(1-2\eta_{l})K,$ (101)
which means that $\Delta$, the gap of Hamiltonian $K$, satisfies
$\Delta\geq 1-2\eta_{l}.$ (102)
To prove that the above bound is included in the feasible set, we find a positive and local decomposition for $Q_{3l}(1-2\eta_{l})=[K^{2}]_{3l}-(1-2\eta_{l})K$, where the subscript $3l$ refers to truncation size in the original system
$Q_{3l}(1-2\eta)$ $=[K^{2}]_{3l}-(1-2\eta_{l})K$ (103)
$:=\sum_{i}q_{3l,i}(1-2\eta_{l})$ (104)
such that $q_{3l,i}\succeq 0$. We start by expanding the LHS,
$Q_{3l}(1-2\eta_{l})$ $=K+\sum_{\mathcal{I}}\{P_{\mathcal{I}},P_{\mathcal{I}+1}\}-(1-2\eta_{l})K$ (105)
$=\sum_{i}2\eta_{l}P_{\mathcal{I}}+\{P_{\mathcal{I}},P_{\mathcal{I}+1}\}$ (106)
$=\sum_{i}\eta_{l}(P_{\mathcal{I}}+P_{\mathcal{I}+1})+\{P_{\mathcal{I}},P_{\mathcal{I}+1}\}$ (107)
@@
$q_{3l,i}(1-2\eta_{l})\succeq 0.$ (109)
Appendix C Optimizing the interaction term using the SDP gradient formula
@@
First of all, we see that upon applying the SDP gradient formula ([52], Section 4.2) to the exact SDP formulation of the gap problem section 2, we recover the result of the Hellmann–Feynman theorem: Let $H$ be the full system Hamiltonian with ground-state energy zero, and let $P_{0}$ be the projector to the global ground state. Assume $H$ depends on a parameter $x$ as $H(x)=H_{0}+(\mathbb{I}-P_{0})A(x)(\mathbb{I}-P_{0})$, where $A$ is some operator depending smoothly on $x$, such that $H(x)$ has the same ground state as $H_{0}$ in a neighborhood of $x_{0}$.
The SDP gradient formula tells us that the derivative of the objective function (in our case the exact gap) with respect to the parameter $x$ is given by the partial derivative of the Lagrangian at the primal-dual optimal point. In the same spirit as the Hellmann–Feynman theorem—only the explicit dependence on $x$ contributes to the gradient. Applying the SDP gradient formula to the SDP section 2, where the corresponding Lagrangian is $\mathcal{L}(\rho,\delta)=\delta+\mathrm{Tr}(\rho(H^{2}-\delta H))$, and the primal-dual optimum is ($\rho^{\star}=|\psi_{1}\rangle\langle\psi_{1}|/\Delta,\delta^{\star}=\Delta$), we get
@@
Let $p_{0}$ be the projector to the ground space of $h_{0}$. Any interaction term of the form $h(X):=h_{0}+(\mathbb{I}-p_{0})X(\mathbb{I}-p_{0})$ which is still positive semidefinite and frustration free will have the same global ground state. We can use the SDP gradient formula to differentiate the gap which the SDP detects with respect to $X$. Of course we can always increase the gap arbitrarily by choosing $X=\mathrm{const.}\times\mathbb{I}$, so in addition to the positivity constraint $h(X)\succeq 0$ we should impose some constraint on $X$, e.g. $X||_{2}\leq 1$. Both of those constraints are easily differentiable and so we could perform constrained gradient optimization.
@@
Appendix D Certifiable formulation of the LTI SDP
To regard the solution of the SDP as a certificate or proof of a gap $\geq\delta$ it is crucial that we have a handle on numerical precision errors. We now explain how the LTI SDP needs to be modified in order to control for such errors. Recall the LTI SDP: